solve
3a-8=(a/2)+2
a/2 is a fraction, but I didn't know how to write it on here
thanks

Answer by smik(40)

You would start off by bringing all the variables to one side and the constants to another:


3a-8=(a/2)+2


Becomes: 3a-(a/2)=10


Now, we multiply each side by 2 so to get rid of the denominator in a/2:


2(3a-(a/2))=2(10)


Becomes: 6a-a=20


Becomes: 5a=20


Divide each side by 5:


a=4